# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 2.3.2 Pseudosymmetries of differential equations
Then by $\left[X_{s}(\xi^{h})\right]\left[\nu_{h}^{j}\right]=\left[X_{s}(\nu^{j})\right]$ one readily gets $\left[X_{i}(\xi^{h})\right]\left[Y(\nu_{h}^{j})\right]=0$. The result follows from non-degeneracy condition $\det\left[X_{i}(\xi^{h})\right]\neq 0$. An analogous proof holds in the case of an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$, since for any $X_{i}$ and $Y_{h}$ one has $[Y_{h},X_{i}]=\sum_{s}\alpha_{hi}^{s}X_{s}+\sum_{k}\beta_{hi}^{k}Y_{k}$, for some functions $a_{hi}^{s}$ and $\beta_{hi}^{k}$. ∎
In general one has no guarantee that the “derived” system of invariants $\{\xi^{i},\nu^{j},\nu_{i}^{j}\}$ is functionally independent, even when $\{\xi^{i},\nu^{j}\}$ is a functionally independent system. The scenario, however, changes in the case of the Cartan distribution $\mathcal{C}(\pi)$ on the infinite jet space. Indeed, in such a case the generators $X_{i}$ of $\mathcal{C}(\pi)$ are the total derivatives, hence the jet-degree of $\left[\nu_{h}^{j}\right]=\left[X_{i}(\xi^{h})\right]^{-1}\left[X_{i}(\nu^{j})\right]$ (which is (15)) is higher than that of $\{\xi^{i},\nu^{j}\}$. Thus, if $\left\{\xi^{i},\nu^{j}\right\}$ are functionally independent invariants of a pseudosymmetry $Y$ (or, an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$) of $\mathcal{\mathcal{C}}(\pi)=<D_{1},...,D_{n}>$ satisfying $\det\left(D_{s}\xi^{i}\right)\neq 0$, the prolongation formulas (5) provide an infinite sequence $\{x_{i}^{\prime}=\xi^{i},\,u^{\prime}{}_{\mu}^{j}=\nu_{\mu}^{j}\}$ of invariants of $Y$ (resp., $\{Y_{1},...,Y_{r}\}$) that defines a regular $\mathcal{C}$-morphism in $J^{\infty}(\pi)$, in view of Proposition 3.
On the other hand, given a formally integrable equation =$\mathcal{E}=\{\mathbf{F}=0\}\subset J^{k}(\pi)$, if $\left\{\xi^{i},\nu^{j}\right\}$ are functionally independent invariants of a pseudosymmetry $\bar{Y}$ (or, an $r$-pseudosymmetry $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$) of $\mathcal{\mathcal{C}}(\mathcal{E})=<\bar{D}_{1},...,\bar{D}_{n}>$ on $\mathcal{E}^{(\infty)}$ satisfying $\det\left(\bar{D}_{s}\xi^{i}\right)\neq 0$, then prolongation formulas (15) allow one to define a sequence $\{\xi^{i},\,\nu_{\mu}^{j}\}$ of invariants of $\bar{Y}$ (resp., $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$) that are functionally dependent, due to the constraints imposed by $\mathbf{F}=0$ and its differential consequences. Indeed, assuming for instance that $\{\xi^{i},\,\nu_{\mu}^{j}\}$ are $\ell$-th order invariants of $\bar{Y}$, the cardinality of $\{\xi^{i},\,\nu_{\mu}^{j}\}$ grows with $|\mu|$ exactly as the dimension of $J^{|\mu|}(\pi)$, that in its turn grows faster than the dimension of the prolongation $\mathcal{E}^{(\ell+|\mu|-k)}$. Therefore, for sufficiently large $|\mu|$ the system $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is necessarily functionally dependent, in view of the fact that for a nontrivial $\bar{Y}$ there are at most $\dim\mathcal{E}^{(\ell+|\mu|-k)}-1$ functionally independent invariants of order $\ell+|\mu|$, as one can see by considering the projections of $\bar{Y}$ on the prolongations of $\mathcal{E}$. Indeed, looking at vector fields as derivations, the projection $\left.\bar{Y}\right|_{\mathcal{E}}$ of $\bar{Y}$ on $\mathcal{E}$ is defined as the restriction of $\bar{Y}$ to the algebra of smooth functions on $\mathcal{E}$. Analogously, for $\mathcal{q}>k$, the projection $\left.\bar{Y}\right|_{\mathcal{E}^{(\mathcal{q}-k)}}$ of $\bar{Y}$ on $\mathcal{E}^{(\mathcal{q}-k)}$ is defined as the restriction of $\bar{Y}$ to the algebra of smooth functions on $\mathcal{E}^{(\mathcal{q}-k)}$. When the projection of $\bar{Y}$ to $\mathcal{E}^{(\mathcal{q}-k)}$ (or $\mathcal{E}$) is a derivation, hence a vector field on $\mathcal{E}^{(\mathcal{q}-k)}$ (or $\mathcal{E}$, resp.), one has that invariants of the projection are also invariants of $\bar{Y}$. On the contrary, when the projection of $\bar{Y}$ to $\mathcal{E}^{(\mathcal{q}-k)}$ (or $\mathcal{E}$) is not a derivation, hence it is a relative vector field on $\mathcal{E}^{(\mathcal{q}-k)}$ (or $\mathcal{E}$, resp.), one has that invariants of the projection are not necessarily also invariants of $\bar{Y}$. Thus, in any case, the number of functionally independent invariants of order $\mathcal{q}$ (or $k$, resp.) admitted by $\bar{Y}$ is at most $\dim(\mathcal{E}^{(\mathcal{q}-k)})-1$ (or $\dim(\mathcal{E})-1$, resp.). Analogous considerations hold when $\{\xi^{i},\,\nu_{\mu}^{j}\}$ are invariants of an $r$-pseudosymmetry.
Therefore, since $\mathbf{F}=0$ and its differential consequences entail that for sufficiently large $|\mu|$ the system $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is functionally dependent on $\mathcal{E}^{(\infty)}$, it follows that the invariants $\{\nu{}_{\sigma}^{j}\}$ satisfy an associated system of partial differential equations $\mathbf{Q}=0$ (together with their prolongations), modulo $\mathbf{F}=0$ and its differential consequences. This way one can look at $\{x^{\prime}_{i}=\xi^{i},\,u^{\prime}{}_{\mu}^{j}=\nu_{\mu}^{j}\}$ as the components of a $\mathcal{C}$-morphism from =$\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}$ to $\mathcal{Q}=\{\mathbf{Q}=\mathbf{0}\}$. Of course, if one chooses another system of invariants $\{\bar{x}^{\prime}_{i}=\bar{\xi}^{i},\,\bar{u}^{\prime}{}^{j}=\bar{\nu}^{j}\}$ equivalent to $\{\xi^{i},\,\nu^{j}\}$, i.e., such that $\{\bar{\xi}^{i}=A^{i}(\xi^{i},\nu^{j}),\,\bar{\nu}^{j}=B^{j}(\xi^{i},\nu^{j})\}$ and $\det\left(D_{s}\bar{\xi}^{i}\right)\neq 0$, one would obtain a system $\mathcal{Q}^{\prime}$ which is point equivalent to $\mathcal{Q}$, through the point transformation $\{\bar{x}^{\prime}_{i}=A^{i}(x^{\prime},u^{\prime}),\,\bar{u}^{\prime}{}^{j}=B^{j}(x^{\prime},u^{\prime})\}$. In particular, one can choose $\left\{\xi^{i},\nu^{j}\right\}$ in such a way that none of its element is the prolongation of another one and, in addition to that, $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is a complete system of invariants, i.e., a system that generates invariants of all possible orders. In such a case, according to the common terminology used in the particular case of symmetries (see for instance [22, 38, 39]), one can look at $\mathcal{Q}$ as the factorization (or the quotient) of $\mathcal{E}$ by $\bar{Y}$.
Examples 13 and 14 below will provide a simple illustration of above factorization procedure. Other examples, describing Bäcklund transformations, will be discussed in Section 4.
In view of these considerations, we are particularly interested here to the determination of pseudosymmetries or $r$-pseudosymmetries of $\mathcal{\mathcal{C}}(\mathcal{E})$ on $\mathcal{E}^{(\infty)}$. Using the outer geometry of $J^{\infty}(\pi)$, we will consider pseudosymmetries and $r$-pseudosymmetries of $\mathcal{\mathcal{C}}(\mathcal{E})$ that are obtained by restricting to $\mathcal{E}^{(\infty)}$ suitable vector fields on $J^{\infty}(\pi)$. Thus, it is convenient to adopt the following
@@
Let =$\mathcal{E}=\{\mathbf{F=0}\}\subset J^{k}(\pi)$ be a formally integrable differential equation. A pseudosymmetry of $\mathcal{E}$ is a vector field $Y$ on $J^{\infty}(\pi)$ which is tangent to $\mathcal{E}^{(\infty)}$ and restricts to a pseudosymmetry $\bar{Y}:=\left.Y\right|_{\mathcal{E}^{(\infty)}}$ of $\mathcal{C}(\mathcal{E})$. Analogously, an $r$-pseudosymmetry of $\mathcal{E}$ is a system of vector fields $\{Y_{1},...,Y_{r}\}$ on $J^{\infty}(\pi)$ which are tangent to $\mathcal{E}^{(\infty)}$ and restrict to an $r$-pseudosymmetry $\{\bar{Y}_{1}:=\left.Y_{1}\right|_{\mathcal{E}^{(\infty)}},...,\bar{Y}_{r}:=\left.Y_{r}\right|_{\mathcal{E}^{(\infty)}}\}$ of $\mathcal{C}(\mathcal{E})$.
In his paper [38], Sokolov discussed a method for calculating the pseudosymmetries of $\mathcal{E}$. We will describe this method here using the following proposition, along with a short proof for the convenience of our readers.
@@
[…]
$[D_{i}+U_{i},D_{j}+U_{j}]=0\mod{\mathcal{E}^{(\infty)}}.$ (16)
@@
$Y=\sum_{i}a_{i}D_{i}+\sum_{|\sigma|\geq 0}\sum_{j}\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}},$ (17)
@@
is a pseudosymmetry of $\mathcal{E}$ whenever $Y(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, i.e.,
@@
for any $h=1,...,p$, which is equivalent to the tangency condition of $Y$ to $\mathcal{E}^{(\infty)}$.
?proofname?.
@@
$\begin{array}[]{ll}\left[D_{s},Y\right]&=\left[D_{s},a_{i}D_{i}\right]+\left[D_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right]=D_{s}(a_{i})D_{i}+\left[D_{s}+U_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right]-\left[U_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right].\vspace{5pt}\end{array}$
@@
$\begin{array}[]{ll}\left[D_{s},Y\right]&=D_{s}(a_{i})D_{i}+\left(D+U\right)_{\sigma+1_{s}}(\varphi^{j})\partial_{u_{\sigma}^{j}}-\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma-1_{s}}^{j}}-\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\circ U_{s}+\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\circ U_{s}\vspace{5pt}\\ &\qquad-U_{s}\circ(Y-a_{i}D_{i})\mod{\mathcal{E}^{(\infty)}},\end{array}$
@@
$[D_{s},Y]=\left(D_{s}(a_{i})+a_{i}U_{s}\right)D_{i}-U_{s}Y\mod{\mathcal{E}^{(\infty)}},$ (19)
@@
$D_{s}\left(Y(D_{\sigma}(F^{h}))\right)-Y\left(D_{s}(D_{\sigma}(F^{h}))\right)=\left(D_{s}(a_{i})+a_{i}U_{s}\right)D_{i}\left(D_{\sigma}(F^{h})\right)-U_{s}Y(D_{\sigma}(F^{h}))=-U_{s}Y(D_{\sigma}(F^{h}))\mod{\mathcal{E}^{(\infty)}},$
@@
$Y\left(D_{\sigma+1_{s}}(F^{h})\right)=U_{s}Y(D_{\sigma}(F^{h}))+D_{s}\left(Y(D_{\sigma}(F^{h}))\right)\mod{\mathcal{E}^{(\infty)}},$
for any $\sigma$ and for any $h$. Thus $Y(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, $h=1,...,p$, guarantee that $Y\left(D_{\sigma}(F^{h})\right)=0\mod{\mathcal{E}^{(\infty)}}$, i.e., that $Y$ is tangent to $\mathcal{E}^{(\infty)}$. Therefore, in view of (19), one gets that $Y$ restricts on $\mathcal{E}^{(\infty)}$ to a pseudosymmetry of $\mathcal{C}(\mathcal{E})$. ∎
Some remarks are in order here. First, in view of (17), one has recurrence formulas for the components of a pseudosymmetry that allow one to obtain higher order components from lower order ones. To distinguish the case of pseudosymmetries from that of symmetries, we will say that (17) is the pseudo-prolongation, relatively to the conservation law $\mu$, of the (possibly relative) vector field $\sum_{i}a_{i}\partial_{x_{i}}+\sum_{j}\varphi^{j}\partial_{u^{j}}$ on $J^{0}(\pi)$. In particular, when $\mu=0$ the pseudo-prolongation reduces to standard prolongation and the pseudosymmetry is just a symmetry with generating function $\varphi$. Also, it is important to observe that in practice the determination of pseudosymmetries by means of (18) only requires the consideration of a finite number of differential consequences of $\mathcal{E}$.
@@
[…]
$[D_{i}+\mathbb{U}_{i},D_{j}+\mathbb{U}_{j}]=0\mod{\mathcal{E}^{(\infty)}}.$ (20)
@@
where $\mathbb{D}:=(D_{1},...,D_{n})^{T}$, $\partial_{\mathbf{u}_{\sigma}}=(\partial_{u_{\sigma}^{1}},...,\partial_{u_{\sigma}^{m}})^{T}$, $\mathbb{A}$ is an $r\times n$ matrix function, $\Phi$ is an $r\times m$ matrix function and
@@
describes an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$ of $\mathcal{E}$ whenever $\mathbb{Y}(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, i.e.,
@@
for any $h=1,...,p$, which is equivalent to the tangency condition of each $Y_{1},...,Y_{r}$ to $\mathcal{E}^{(\infty)}$.
For $r$-pseudosymmetries apply same considerations made above for pseudosymmetries. In particular, one can say that the system (21) is the pseudo-prolongation, relatively to the ZCR $\gamma$, of the (possibly relative) vector fields $\sum_{i}a_{i}^{h}\partial_{x_{i}}+\sum_{j}\varphi^{hj}\partial_{u^{j}}$, $h=1,...,r$, on $J^{0}(\pi)$.

### v5

@@ 2.3.2 Pseudosymmetries of differential equations and factorisation
Then by $\left[X_{s}(\xi^{h})\right]\left[\nu_{h}^{j}\right]=\left[X_{s}(\nu^{j})\right]$ one readily gets $\left[X_{i}(\xi^{h})\right]\left[Y(\nu_{h}^{j})\right]=0$. The result follows from non-degeneracy condition $\det\left[X_{i}(\xi^{h})\right]\neq 0$. An analogous proof holds in the case of an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$, since for any $X_{i}$ and $Y_{h}$ one has $Y_{h},X_{i}]=\sum_{s}\alpha_{hi}^{s}X_{s}+\sum_{k}\beta_{hi}^{k}Y_{k}$, for some functions $a_{hi}^{s}$ and $\beta_{hi}^{k}$. ∎
Concerning the pseudosymmetries or $r$-pseudosymmetries of $\mathcal{\mathcal{C}}(\mathcal{E})$ on $\mathcal{E}^{(\infty)}$, it is convenient to adopt the following
@@
Let $\mathcal{E}=\{\mathbf{F=0}\}\subset J^{k}(\pi)$ be a formally integrable differential equation. A pseudosymmetry of $\mathcal{E}$ is a vector field $\bar{Y}$ on $\mathcal{E}^{(\infty)}$ that is a pseudosymmetry of $\mathcal{C}(\mathcal{E})$. Analogously, an $r$-pseudosymmetry of $\mathcal{E}$ is a system of vector fields $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$ on $\mathcal{E}^{(\infty)}$ that is an $r$-pseudosymmetry of $\mathcal{C}(\mathcal{E})$.
In his paper [41], Sokolov discussed a method for calculating pseudosymmetries of $\mathcal{E}$. Here we will describe that method by means of the following proposition, along with a short proof for the convenience of our readers.
@@
Let $\mathcal{E}=\{\mathbf{F=0}\}\subset J^{k}(\pi)$, with $\text{$\mathbf{F}$}=(F^{1},...,F^{p})$, be a formally integrable differential equation and $\mu=U_{1}dx^{1}+...+U_{n}dx^{n}$ an horizontally closed $1$-form on $\mathcal{E}^{(\infty)}$, i.e., such that
$D_{i}+U_{i},D_{j}+U_{j}]=0\mod{\mathcal{E}^{(\infty)}}.$ (16)
@@
$Y=\sum_{i}a_{i}\partial_{x_{i}}+\sum_{|\sigma|\geq 0}\sum_{j}b_{\sigma}^{j}\partial_{u_{\sigma}^{j}}=\sum_{i}a_{i}D_{i}+\sum_{|\sigma|\geq 0}\sum_{j}\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}},$ (17)
@@
$\begin{array}[]{l}\varphi^{j}:=Y\,\lrcorner\,\omega_{0}^{j},\qquad\qquad\omega_{0}^{j}=du^{j}-{\displaystyle\sum_{i}}u_{i}^{j}dx_{i},\vskip 10.0pt\\ b_{\sigma}^{j}:={\displaystyle\sum_{i}a_{i}\,u_{\sigma+1_{i}}^{j}+\sum_{|\sigma|\geq 0}\sum_{j}\left(D+U\right)_{\sigma}(\varphi^{j})},\vskip 10.0pt\\ \left(D+U\right)_{\sigma}:=\left(D_{1}+U_{1}\right)^{\sigma_{1}}\circ...\circ\left(D_{n}+U_{n}\right)^{\sigma_{n}},\vskip 10.0pt\end{array}$
restricts on $\mathcal{E}^{(\infty)}$ to a pseudosymmetry $\bar{Y}$ of $\mathcal{E}$ whenever $Y(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, i.e.,
@@
for any h=1,,ph=1,...,p, which is equivalent to the tangency condition of $Y$ to $\mathcal{E}^{(\infty)}$.
Proof.
@@
$\begin{array}[]{ll}\left[D_{s},Y\right]&=\left[D_{s},a_{i}D_{i}\right]+\left[D_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right]=D_{s}\left(a_{i}\right)\,D_{i}+\left[D_{s}+U_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right]-\left[U_{s},\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\right].\vskip 5.0pt\end{array}$
@@
$\begin{array}[]{ll}\left[D_{s},Y\right]&=D_{s}(a_{i})D_{i}+\left(D+U\right)_{\sigma+1_{s}}(\varphi^{j})\partial_{u_{\sigma}^{j}}-\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma-1_{s}}^{j}}-\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\circ U_{s}+\left(D+U\right)_{\sigma}(\varphi^{j})\partial_{u_{\sigma}^{j}}\circ U_{s}\vskip 5.0pt\\ &\qquad-U_{s}\circ(Y-a_{i}D_{i})\mod{\mathcal{E}^{(\infty)}},\end{array}$
@@
$D_{s},Y]=\left(D_{s}a_{i}+a_{i}U_{s}\right)D_{i}-U_{s}Y\mod{\mathcal{E}^{(\infty)}},$ (19)
@@
$D_{s}\left(Y(D_{\sigma}F^{h})\right)-Y\left(D_{s}(D_{\sigma}F^{h})\right)=\left(D_{s}a_{i}+a_{i}U_{s}\right)\,D_{i}\left(D_{\sigma}F^{h}\right)-U_{s}\,Y(D_{\sigma}F^{h})=-U_{s}\,Y(D_{\sigma}F^{h})\mod{\mathcal{E}^{(\infty)}},$
@@
$Y\left(D_{\sigma+1_{s}}F^{h}\right)=U_{s}\,Y(D_{\sigma}F^{h})+D_{s}\left(Y(D_{\sigma}F^{h})\right)\mod{\mathcal{E}^{(\infty)}},$ (20)
for any $\sigma$ and for any $h$. Thus, by repeatedly using (20) with increasing $|\sigma|$, conditions $Y(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, h=1,,ph=1,...,p, guarantee that $Y\left(D_{\sigma}F^{h}\right)=0\mod{\mathcal{E}^{(\infty)}}$, i.e., that $Y$ is tangent to $\mathcal{E}^{(\infty)}$. Therefore, in view of (19), one gets that $Y$ restricts on $\mathcal{E}^{(\infty)}$ to a pseudosymmetry $\bar{Y}$ of $\mathcal{C}(\mathcal{E})$. ∎
Some remarks are in order here. First, in view of (17), one has recurrence formulas for the components of a pseudosymmetry that allow one to obtain higher order components from lower order ones. To distinguish the case of pseudosymmetries from that of symmetries, we will say that (17) is the pseudoprolongation, relatively to the $1$-form $\mu$, of the (possibly relative) vector field $\sum_{i}a_{i}\partial_{x_{i}}+\sum_{j}b^{j}\partial_{u^{j}}$ on $J^{0}(\pi)$. In particular, when $\mu=0$ the pseudoprolongation reduces to standard prolongation (11) and $Y$ is a symmetry of $\mathcal{C}(\pi)$; thus $\varphi=(\varphi^{j})$ is the generating function of a (classical or generalised) symmetry of $\mathcal{E}$ whenever $\mu=0$ and $Y$ is tangent to $\mathcal{E}^{(\infty)}$. Also, we remark that for any given $1$-form $\mu$, the system (18) is a linear system of PDEs for $\varphi=(\varphi^{j})$ that involves only a finite number of differential consequences of $\mathcal{E}$, whenever the order of $\varphi$ is finite. As with the standard symmetry condition, the analysis of such a system of linear PDEs is generally more feasible if one uses some symbolic manipulation package of the type developed in a computer algebra system like Maple. Finally, we note that, as with generalised symmetries, also pseudosymmetries could be put in an evolutionary form (see [16, 41]). In this article, however, we will not limit ourselves to considering pseudosymmetries in evolutionary form, because in general the space of invariants of a pseudosymmetry of $\mathcal{E}$ is not the same as that of its evolutionary form, and this is a non-negligible fact when, as in this paper, one is interested in studying factorisation by pseudosymmetries.
@@
Let $\mathcal{E}=\{\mathbf{F=0}\}\subset J^{k}(\pi)$, with $\text{$\mathbf{F}$}=(F^{1},...,F^{p})$, be a formally integrable differential equation and $\gamma=\mathbb{U}_{1}dx^{1}+...+\mathbb{U}_{n}dx^{n}$ a horizontal $1$-form with $r\times r$ matrix-valued components satisfying the zero-curvature condition $d_{H}\gamma-\gamma\wedge\gamma=0\mod{\mathcal{E}^{(\infty)}},$ i.e.,
$D_{i}+\mathbb{U}_{i},D_{j}+\mathbb{U}_{j}]=0\mod{\mathcal{E}^{(\infty)}}.$ (21)
@@
where $\mathbb{D}:=(D_{1},...,D_{n})^{T}$, $\partial_{\mathbf{u}_{\sigma}}=(\partial_{u_{\sigma}^{1}},...,\partial_{u_{\sigma}^{m}})^{T}$, $\mathbb{A}$ is an $r\times n$ matrix function, $\Phi=(\varphi_{i}^{j})$ is an $r\times m$ matrix function with $\varphi_{i}^{j}:=Y_{i}\,\lrcorner\,\omega_{0}^{j}$ and
@@
defines on $\mathcal{E}^{(\infty)}$ an $r$-pseudosymmetry $\{\bar{Y}_{1}=\left.Y_{1}\right|_{\mathcal{E}^{(\infty)}},...,\bar{Y}_{r}=\left.Y_{r}\right|_{\mathcal{E}^{(\infty)}}\}$ of $\mathcal{E}$ whenever $\mathbb{Y}(F^{h})=0\mod{\mathcal{E}^{(\infty)}}$, i.e.,
@@
for any h=1,,ph=1,...,p, which is equivalent to the tangency condition of each $Y_{1},...,Y_{r}$ to $\mathcal{E}^{(\infty)}$.
For $r$-pseudosymmetries apply same considerations made above for pseudosymmetries. In particular, one can say that the system (22) is the pseudoprolongation, relatively to $\gamma$, of the (possibly relative) vector fields $\sum_{i}a_{hi}\partial_{x_{i}}+\sum_{j}b_{h}^{j}\partial_{u^{j}}$, h=1,,rh=1,...,r (with $\varphi_{h}^{j}=b_{h}^{j}-\sum_{i}a_{hi}u_{i}^{j})$ on $J^{0}(\pi)$.
In general one has no guarantee that the system of “derived” (or prolonged) invariants $\{\xi^{i},\nu^{j},\nu_{i}^{j}\}$ (provided by Proposition 9) is functionally independent, even when $\{\xi^{i},\nu^{j}\}$ is a functionally independent system. The scenario, however, changes in the case of the Cartan distribution $\mathcal{C}(\pi)$ on the infinite jet space, for an $m$-dimensional bundle $\pi:E\rightarrow M$ with $n=\dim M$. Indeed, in such a case the generators $X_{i}$ of $\mathcal{C}(\pi)$ are the total derivatives, hence the jet-order of $\left[\nu_{h}^{j}\right]=\left[D_{i}\xi^{h}\right]^{-1}\left[D_{i}\nu^{j}\right]$ (that is equation (15)) is higher than that of $\{\xi^{i},\nu^{j}\}$. Thus, if $\left\{\xi^{i},\nu^{j}\right\}$ are functionally independent invariants of a pseudosymmetry $Y$ (or, an $r$-pseudosymmetry $\{Y_{1},...,Y_{r}\}$) of $\mathcal{\mathcal{C}}(\pi)=<D_{1},...,D_{n}>$ satisfying $\det\left(D_{s}\xi^{i}\right)\neq 0$, the prolongation formulas (15) provide an infinite sequence $\{x_{i}^{\prime}=\xi^{i},\,u^{\prime}{}_{\mu}^{j}=\nu_{\mu}^{j}\}$ of invariants of $Y$ (resp., $\{Y_{1},...,Y_{r}\}$); when $p=m$ this sequence defines a regular $\mathcal{C}$-morphism in $J^{\infty}(\pi)$, in view of Proposition 3.
On the other hand, given a formally integrable equation $\mathcal{E}=\{\mathbf{F}=0\}\subset J^{k}(\pi)$, if $\left\{\xi^{1},...,\xi^{n},\nu^{1},...,\nu^{p}\right\}$, $p\geq m$, are functionally independent invariants of a pseudosymmetry $\bar{Y}$ (resp., $r$-pseudosymmetry $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$) of $\mathcal{\mathcal{C}}(\mathcal{E})=<\bar{D}_{1},...,\bar{D}_{n}>$ on $\mathcal{E}^{(\infty)}$ satisfying $\det\left(\bar{D}_{s}\xi^{i}\right)\neq 0$, then prolongation formulas (15) allow one to define a sequence $\{\xi^{i},\,\nu_{\mu}^{j}\}$ of invariants of $\bar{Y}$ (resp., $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$) that are functionally dependent, due to the constraints imposed by $\mathbf{F}=0$ and its differential consequences. Indeed, assuming for instance that $\{\xi^{i},\,\nu_{\mu}^{j}\}$ are $\ell$-th order invariants of $\bar{Y}$, the cardinality of $\{\xi^{i},\,\nu_{\mu}^{j}\}$ grows with $|\mu|$ faster than the dimension of the prolongation $\mathcal{E}^{(\ell+|\mu|-k)}$. Therefore, for sufficiently large $|\mu|$ the system $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is necessarily functionally dependent. In particular, a nontrivial $\bar{Y}$ has at most $\dim\mathcal{E}^{(\ell+|\mu|-k)}-1$ functionally independent invariants of order $\ell+|\mu|\geq k$, with the maximum reached only when $\bar{Y}$ projects (or restricts) to a nontrivial vector field (i.e., a derivation) on $\mathcal{E}^{(\ell+|\mu|-k)}$. Indeed, when $\bar{Y}$ projects to a relative vector field on $\mathcal{E}^{(\ell+|\mu|-k)}$, the invariants of the projection are not necessarily also invariants of $\bar{Y}$. Analogous considerations hold when $\{\xi^{i},\,\nu_{\mu}^{j}\}$ are invariants of an $r$-pseudosymmetry.
Therefore, since $\mathbf{F}=0$ and its differential consequences entail that for sufficiently large $|\mu|$ the system $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is functionally dependent on $\mathcal{E}^{(\infty)}$, it follows that the invariants $\{\nu{}_{\sigma}^{j}\}$ satisfy an associated system of partial differential equations $\mathbf{Q}=0$ (together with their prolongations), modulo $\mathbf{F}=0$ and its differential consequences. This way one can look at $\{x^{\prime}_{i}=\xi^{i},\,u^{\prime}{}_{\mu}^{j}=\nu_{\mu}^{j}\}$ as the components of a $\mathcal{C}$-morphism from $\mathcal{E}=\{\mathbf{F}=\mathbf{0}\}$ to $\mathcal{Q}=\{\mathbf{Q}=\mathbf{0}\}$. Of course, if one chooses another system of invariants $\{\hat{x}^{\prime}_{i}=\hat{\xi}^{i},\,\hat{u}^{\prime}{}^{j}=\hat{\nu}^{j}\}$ equivalent to $\{\xi^{i},\,\nu^{j}\}$, i.e., such that $\{\hat{\xi}^{i}=A^{i}(\xi^{i},\nu^{j}),\,\hat{\nu}^{j}=B^{j}(\xi^{i},\nu^{j})\}$ and $\det\left(\bar{D}_{s}\hat{\xi}^{i}\right)\neq 0$ (where $\bar{D}_{s}$ are the total derivatives on $\mathcal{E}$), one would obtain a system $\mathcal{\hat{Q}}$ which is point equivalent to $\mathcal{Q}$, through the point transformation $\{\hat{x}^{\prime}_{i}=A^{i}(x^{\prime},u^{\prime}),\,\hat{u}^{\prime}{}^{j}=B^{j}(x^{\prime},u^{\prime})\}$. In particular, one can choose $\left\{\xi^{i},\nu^{j}\right\}$ in such a way that none of its element is the prolongation of another one and, in addition to that, $\{\xi^{i},\,\nu_{\mu}^{j}\}$ is a complete system of invariants, i.e., a system that generates invariants of all possible orders. In such a case, according to the common terminology used in the particular case of symmetries (see for instance [25, 41, 42]), one can look at $\mathcal{Q}$ as the factorisation, or the quotient, of $\mathcal{E}$ by $\bar{Y}$ (resp., $\{\bar{Y}_{1},...,\bar{Y}_{r}\}$).
Next two examples will provide a simple illustration of above factorisation procedure. Other examples, describing Bäcklund transformations, will be discussed in Section 4.